How old are they?
Jack's dad is 40. Jack is 14. How old will each of them be when his dad is twice as old as Jack?

Answer by checkley75(3666)

40+X=2(14+X)
40+X=28+2X
X-2X=28-40
-X=-12
X=12 YEARS FROM NOW DAD WILL BE
40+12=52 &
JACK WILL BE 14+12=26
THUS IN 12 YEARS DAD WILL BE TWICE AS OLD AS JACK
2*26=52